OpenStudy (anonymous):

A cylindrical oil tank with a radius of 2 meters holds 1,000 cubic meters of oil. What is the approximate height of the tank? Use 3.14 for pi, and round your answer to the nearest meter.

OpenStudy (nurali):

area of base: pi(r^2) = 3.14(2^2) =12.28 m^2. Volume = h x area of base = h x 12.28 = 1000 h = 1000/12.28 = 81.4 meters, rounds to 81 meters.